OpenStudy (anonymous):

Find the second derivative of f(x)=x^(5/3)+2 and show steps.

OpenStudy (anonymous):

have you found the first derivaitve

OpenStudy (anonymous):

I think so. I got (5/3)x^(2/3) which I don't know if it is correct but I've had problems with the chain rule and don't know if there was more I was supposed to do..

OpenStudy (anonymous):

It looks like a basic power rule problem: f(x)=x^(5/3)+2 f '(x)=(5/3)x^(2/3) Then take the derivative again.

OpenStudy (anonymous):

Awesome so I am on the right track!

OpenStudy (anonymous):

f'(x)=(5/3)x^(2/3) f"(x)=(10/9)x^(-1/3)

OpenStudy (anonymous):

So this is my first time posting. Once this is answered should I delete it or leave it? Whats the proper etiquette?

OpenStudy (amistre64):

questions asked and answered are for the benefit of the community as a whole; so its best to leave it up. Deleting is for when you post bad to begin with.

OpenStudy (anonymous):

you dont wanna erase my work do you :(

OpenStudy (anonymous):

Haha I'd hate to erase this great work, thanks!